OpenStudy (anonymous):

-1/2 + 3 -------- -8 is this right and where do i go next

OpenStudy (anonymous):

you can still simplify the top, which is 3 -.5

OpenStudy (anonymous):

so once you do that you have the 2.5 on top and -8 on the bottom

OpenStudy (anonymous):

the 2.5/-8 is the final answer for the slope and is in fraction form

OpenStudy (anonymous):

first add the numerator u get2.5 2.5=2,1,2 then divide

OpenStudy (anonymous):

= 5/(2*-8) = - 5/16 You can go wherever you want,ask answer, welcome aboard.

OpenStudy (anonymous):

Are we all talking about same thing here? I see (-1/2 +3)/-8